How many molecules (not moles) of NH3 are produced from 8.45×10−4 g of H2?
1
Expert's answer
2020-02-26T05:01:47-0500

Solution.

"N2 + 3H2 = 2NH3"

"n(H2) = \frac{8.45 \times 10^{-4}}{2} = 4.23 \times 10^{-4} \ mol"

"n(NH3) = \frac{4.23 \times 10^{-4} \times 2}{3} = 2.82 \times 10^{-4} \ mol"

"N(NH3) = n(NH3) \times Na = 2.82 \times 10^{-4} \times 6.02 \times 10^{23} = 1.7 \times 10^{20} \ molecules"

Answer:

"N(NH3) = 1.7 \times 10^{20} \ molecules"